The firm fixed costs is RM100,00 per year. Management expects to sell 2,000 units per year and at that rate of output, TVC will be RM50,000. The firm uses cost-plus pricing to earn a target rate of return on its RM200,000 investment. If the price is set at RM100, what is the target rate of return
1
Expert's answer
2016-12-14T11:32:08-0500
TFC = 100,00*2,000 = 200,000,00
TC = TVC + TFC = 200,000 + 50,000 = 250,000
If price = 100,00, then the real target rate of return is TC + Price*Q = 250,000 + 100,00*2,000 = 450,000.